# Tighter Confidence Bounds for Sequential Kernel Regression

### v1

@@ 5.5 Kernel AMM-UCB
For our third and final variant of KernelUCB, we set $u_{t}(x)$ to be the analytic UCB from Cor. 4.6 with a fixed value of $\alpha$. The resulting algorithm, which we call Kernel Analytic Martingale Mixture UCB (Kernel AMM-UCB), is a kernelised version of the AMM-UCB algorithm for linear bandits (Flynn et al., 2023). This algorithm is mainly useful for our regret analysis in Section 6.2.
@@ 6 Theoretical Analysis
In this section, we show that our analytic confidence bound from Cor. 4.6 is always tighter than the confidence bounds in Thm. 3.11 of (Abbasi-Yadkori, 2012) and Thm. 2 of (Chowdhury & Gopalan, 2017), which hold under the same conditions and assumptions. In addition, we establish cumulative regret bounds for Kernel CMM-UCB, Kernel DMM-UCB and Kernel AMM-UCB. We begin by formally stating the assumptions under which our analysis holds. We assume that the noise variables are conditionally $\sigma$-sub-Gaussian.
Assumption 6.1 (Sub-Gaussian noise).
Let $(\mathcal{D}_{t}|t\in\mathbb{N})$ be any filtration such that $y_{t}$ and $x_{t+1}$ are both $\mathcal{D}_{t}$-measurable. Each noise variable $\epsilon_{t}$ is conditionally $\sigma$-sub-Gaussian, which means
$\forall\lambda\in\mathbb{R},\;\ \mathbb{E}\left[\mathrm{exp}(\lambda\epsilon_{t})|\mathcal{D}_{t-1}\right]\leq\mathrm{exp}(\lambda^{2}\sigma^{2}/2).$
This condition implies that $\mathbb{E}\left[\epsilon_{t}|\mathcal{D}_{t-1}\right]=0$. We assume that the reward function is uniformly bounded, and that its RKHS norm is bounded.
Assumption 6.2 (Bounded reward function).
For some $B>0$, $\left\|f^{*}\right\|_{\mathcal{H}}\leq B$. For some $C>0$, $\sup_{x\in\mathcal{X}}|f^{*}(x)|\leq C$.
If the kernel function is uniformly bounded by some $\kappa>0$, i.e. $k(x,x^{\prime})\in[-\kappa,\kappa]$ for all $x,x^{\prime}\in\mathcal{X}$, then $\left\|f^{*}\right\|_{\mathcal{H}}\leq B$ implies that $f^{*}(x)\in[-C,C]$ holds with $C\leq\kappa B$. To compute our confidence bounds and run our bandit algorithms, we do not need to know the value of $C$, but we do need to know (upper bounds on) $\sigma$ and $B$.
6.1 Tighter Confidence Bounds
The confidence bounds in (Abbasi-Yadkori, 2012) and (Chowdhury & Gopalan, 2017) are the same as our analytic UCB in (14), except that the (scaled) radius $\widetilde{R}_{\alpha,t}/\sqrt{\alpha}$ is replaced with a different quantity. For any values of the regularisation parameters $\lambda>0$ and $\eta>0$ used in (Abbasi-Yadkori, 2012) and (Chowdhury & Gopalan, 2017) respectively, we show that we can set $c$ and $\alpha$ such that $\widetilde{R}_{\alpha,t}/\sqrt{\alpha}$ is strictly less than these other radius quantities. This result is stated in Thm. 6.3 and proved in App. C.2.
Theorem 6.3.
For every $\delta\in(0,1]$ and every value of the regularisation parameter $\lambda>0$, the RHS of (14) with $c=\sigma^{2}/\lambda$ and $\alpha=\lambda$ is strictly less than the UCB in Thm. 3.11 of (Abbasi-Yadkori, 2012).
For every $\delta\in(0,1]$ and every value of the regularisation parameter $\eta>0$, the RHS of (14) with $c=\sigma^{2}/(1+\eta)$ and $\alpha=1+\eta$ is strictly less than the UCB in Thm. 2 of (Chowdhury & Gopalan, 2017).
6.2 Cumulative Regret Bounds
We bound cumulative regret in terms of a kernel-dependent quantity called the maximum information gain, which is defined as
$\gamma_{t}(\alpha):=\max_{x_{1},\dots,x_{t}}\left\{\frac{1}{2}\ln\left(\det\left(\frac{1}{\alpha}\boldsymbol{K}_{t}+\boldsymbol{I}\right)\right)\right\}.$
First, we need to upper bound our radius quantity $\widetilde{R}_{\alpha,t}$ in terms of the maximum information gain. To do so, we show that when $\alpha=\sigma^{2}/c$, the quadratic terms in $\widetilde{R}_{\alpha,t}$ cancel out (see Lemma C.2), and we are left with
$\widetilde{R}_{\sigma^{2}/c,t}$ $\leq\sqrt{2\sigma^{2}\gamma_{t}(\sigma^{2}/c)+\tfrac{\sigma^{2}B^{2}}{c}+2\sigma^{2}\ln(\tfrac{1}{\delta})}.$
We now proceed to upper bound the cumulative regret of Kernel AMM-UCB with $\alpha=\sigma^{2}/c$. Using a standard argument (see e.g. Prop. 1 in (Russo & Van Roy, 2013) and also Lemma D.1 in App. D.1), the regret of any UCB algorithm at round $t$ is (with high probability) upper bounded by the difference between the upper and lower confidence bounds (at $x_{t}$) that it uses. For Kernel AMM-UCB, this bound is
$r_{t}\leq\frac{2\widetilde{R}_{\sigma^{2}/c,t-1}}{\sqrt{\sigma^{2}/c}}\rho_{\sigma^{2}/c,t-1}(x_{t}).$ (15)
From here, our cumulative regret bound in Thm. 6.4 follows from standard arguments, so we defer the proof to App. D.2.
Theorem 6.4.
$r_{1:T}=\mathcal{O}\left(\sqrt{T\gamma_{T}(\tfrac{\sigma^{2}}{c})\left(\gamma_{T}(\tfrac{\sigma^{2}}{c})+\tfrac{1}{c}B^{2}+\ln(\tfrac{1}{\delta})\right)}\right).$
Because the confidence bounds used by Kernel CMM-UCB (with the same $c$) and Kernel DMM-UCB (with the same $c$ and with $\alpha=\sigma^{2}/c\in A$) are never looser than the confidence bounds used by Kernel AMM-UCB, these algorithms also satisfy the bound in (15), which means the cumulative regret bound for Kernel AMM-UCB also applies to Kernel CMM-UCB and Kernel DMM-UCB.
The regret bound in Thm. 6.4 matches the best existing regret bounds for bandit algorithms that follow the KernelUCB recipe in Algorithm 1 (Srinivas et al., 2010; Chowdhury & Gopalan, 2017; Durand et al., 2018; Whitehouse et al., 2023), though its dependence on the maximum information gain is sub-optimal by a factor of $\sqrt{\gamma_{T}}$. However, due to the lower bound in (Lattimore, 2023), this cannot be improved by using a better confidence sequence for $f^{*}$.
For the frequently used RBF and Matérn kernels, the dependence of the maximum information gain on $T$ can be made explicit. For the RBF kernel and with $c\propto 1$, the maximum information gain is $\mathcal{O}((\ln(T))^{d+1})$ (see e.g. (Srinivas et al., 2010; Vakili et al., 2021)), which means the regret bound in Thm. 6.4 is $\mathcal{O}(\sqrt{T}(\ln(T))^{d+1})$. For the Matérn kernel with smoothness parameter $\nu$ and $c\propto T^{-\frac{d}{2d+2\nu}}$, the maximum information gain is $\mathcal{O}(T^{\frac{d}{2d+2\nu}}(\ln(T))^{\frac{2\nu}{2d+2\nu}})$ (see e.g. (Vakili et al., 2021; Whitehouse et al., 2023)), which means the regret bound is $\mathcal{O}(T^{\frac{2d+\nu}{2d+2\nu}}(\ln(T))^{\frac{2\nu}{d+2\nu}})$.
@@ 7 Experiments
We aim to verify that when our confidence bounds replace existing ones, KernelUCB shows better empirical performance at a similar computational cost.
Compared Methods.
We compare: a) Kernel CMM-UCB: cf. Sec. 5.3; b) Kernel DMM-UCB: cf. Sec. 5.4; c) Kernel AMM-UCB: cf. Sec. 5.5; d) AY-GP-UCB: Algorithm 1 with the confidence bounds from Thm. 3.11 of (Abbasi-Yadkori, 2012); e) IGP-UCB: Algorithm 1 with the confidence bounds from Thm. 2 of (Chowdhury & Gopalan, 2017); e) Random: Actions are chosen uniformly at random.
Experimental Setup.
We run each algorithm in several synthetic kernel bandit problems. To generate each reward function, we sample 20 inducing points $z_{1},\dots,z_{20}$ uniformly at random from the hypercube $[0,1]^{d}$ and a 20-dimensional weight vector $\boldsymbol{w}\in\mathbb{R}^{20}$ from a standard normal distribution. The test function is $f^{*}(x)=b\sum_{i=1}^{20}w_{i}k(x,z_{i})$, where the scaling factor $b$ is set such that $\left\|f^{*}\right\|_{\mathcal{H}}=B$. The noise variables $\epsilon_{1},\epsilon_{2},\dots$ are drawn independently from a normal distribution with mean 0 and standard deviation $\sigma$. In each round $t$, the action set $\mathcal{X}_{t}$ consists of 100 $d$-dimensional vectors drawn uniformly at random from the hypercube $[0,1]^{d}$. We present results for $d\in\{2,3,4\}$ and with the RBF and Matérn kernels with length scale $\ell\in\{0.2,0.5\}$. We record the cumulative regret and wall-clock time-per-step over $T=1000$ rounds.
We run each algorithm with $\delta=0.01$ and the true values $B=10$ and $\sigma=0.1$. We run IGP-UCB with the recommend value of $\eta=2/T$. For each of our methods, we set the covariance scale to $c=1$ for the RBF kernel and $c=T^{-\frac{d}{2d+2\nu}}$ for the Matérn kernel. Equivalently, for AY-GP-UCB, we set $\lambda=\sigma^{2}$ for the RBF kernel and $\lambda=\sigma^{2}T^{\frac{d}{2d+2\nu}}$ for the Matérn kernel. We run Kernel AMM-UCB with $\alpha=\sigma^{2}/c$ and Kernel DMM-UCB with $A=\{0.1\sigma^{2}/c,0.3\sigma^{2}/c,\sigma^{2}/c,3\sigma^{2}/c,10\sigma^{2}/c\}$. Due to prohibitive cost, we only run Kernel CMM-UCB for 200 rounds. All experiments were run on a single computer with an Intel i5-1145G7 CPU.
Results.
Fig. 2 shows the cumulative regret of each algorithm with $d=3$, every kernel and every length-scale (results with $d\in\{2,4\}$ are shown in App. F.2). The cumulative regret at $T=1000$ is displayed in Tab. 1. Our Kernel CMM-UCB algorithm consistently achieves the lowest cumulative regret over the first 200 steps. Our Kernel DMM-UCB algorithm closely matches Kernel CMM-UCB over the first 200 steps and consistently achieves the lowest cumulative regret over 1000 steps. Fig. 4 (in App. F.1) shows the wall-clock time per step for each method. IGP-UCB, AY-GP-UCB and Kernel AMM-UCB take approximately the same time per step. As we would expect, since we used a grid $A$ of size 5, the time per step for Kernel DMM-UCB is a little under 5 times that of Kernel AMM-UCB. The time per step for Kernel CMM-UCB grows at the fastest rate in $t$. At $t=200$, the time per step for Kernel CMM-UCB is more than 15 seconds, whereas the time per step for every other method is below 0.02 seconds.
@@ 8 Conclusion
In this paper, we developed new confidence sequences and confidence bounds for sequential kernel regression. In our theoretical analysis, we proved that our new confidence bounds are always tighter than comparable existing confidence bounds. We presented three variants of the KernelUCB algorithm for kernel bandits, which each use different confidence bounds that can be derived from our confidence sequences. The results of our experiments suggest that the Kernel DMM-UCB algorithm should be the preferred choice for practical uses. It closely matches the empirical performance of our effective but expensive Kernel CMM-UCB algorithm, yet it has roughly the same computational cost as our Kernel AMM-UCB method (as well as AY-GP-UCB and IGP-UCB). Together, our theoretical analysis and our experiments suggest that by replacing existing confidence bounds in KernelUCB with our new confidence bounds, we obtain kernel bandit algorithms with better empirical performance, matching performance guarantees and comparable computational cost.
Though in this paper we applied our confidence bounds only to the kernel bandit problem, and the KernelUCB algorithm in particular, they are a generic tool that could be used anywhere that existing confidence bounds are used. For the future, it remains to apply our confidence bounds to other kernel bandit algorithms and other sequential kernelised learning and/or decision-making problems. For instance, one could use our confidence bounds in the $\pi$-GP-UCB algorithm (Janz et al., 2020), which has recently been shown to achieve optimal order regret for the Matérn kernel (Vakili & Olkhovskaya, 2023). In addition, one could apply our confidence bounds to kernelised reinforcement learning (Yang et al., 2020; Liu & Su, 2022; Vakili & Olkhovskaya, 2023) or adaptive control (Kakade et al., 2020).
Finally, we comment on some limitations of our work. We assume that the kernel as well as reasonably good upper bounds on $\sigma$ and $B$ are known in advance, which is often not the case in practice. While using a loose upper bound on $\sigma$ appears to have little impact, we found that using a loose bound on $B$ can harm empirical performance (see Fig. 7). In future work, we would like to address this limitation by using model selection methods to learn an upper bound on $\left\|f^{*}\right\|_{\mathcal{H}}$, and the kernel or its hyperparameters, from data as it becomes available.
Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v2

@@ 5.5 Kernel AMM-UCB
For our third variant of KernelUCB, we set $u_{t}(x)$ to be the analytic UCB from Cor. 4.6 with a fixed value of $\alpha$. The resulting algorithm is called Kernel Analytic Martingale Mixture UCB (Kernel AMM-UCB).
@@ 6 THEORETICAL ANALYSIS
In this section, we show that our analytic confidence bound from Cor. 4.6 is always tighter than the confidence bounds in Thm. 3.11 of (Abbasi-Yadkori,, 2012) and Thm. 2 of (Chowdhury and Gopalan,, 2017), which hold under the same conditions and assumptions. In addition, we establish cumulative regret bounds for our KernelUCB algorithms. We begin by formally stating the assumptions under which our analysis holds. We assume that $\mathcal{X}$ is a compact subset of $\mathbb{R}^{d}$. In the noisy setting, we assume that the noise variables are conditionally $\sigma$-sub-Gaussian, which means
$\forall\lambda\in\mathbb{R},\;\ \mathbb{E}\left[\mathrm{exp}(\lambda\epsilon_{t})|\mathcal{D}_{t-1}\right]\leq\mathrm{exp}(\lambda^{2}\sigma^{2}/2).$ (15)
For some $B,C>0$, we assume that $\left\|f^{\star}\right\|_{\mathcal{H}}\leq B$ and $\sup_{x}|f^{\star}(x)|\leq C$. If the kernel function is uniformly bounded by some $\kappa>0$, i.e. $\sup_{x,x^{\prime}}|k(x,x^{\prime})|\leq\kappa$, then $\left\|f^{\star}\right\|_{\mathcal{H}}\leq B$ implies that $C\leq\kappa B$.
6.1 Tighter Confidence Bounds
The confidence bounds in (Abbasi-Yadkori,, 2012) and (Chowdhury and Gopalan,, 2017) are the same as our analytic UCB in (14), except that the (scaled) radius $\widetilde{R}_{\alpha,t}/\sqrt{\alpha}$ is replaced with a different quantity. For any values of the regularisation parameters $\lambda>0$ and $\eta>0$ used in (Abbasi-Yadkori,, 2012) and (Chowdhury and Gopalan,, 2017) respectively, we show that we can set $c$ and $\alpha$ such that $\widetilde{R}_{\alpha,t}/\sqrt{\alpha}$ is strictly less than these other radius quantities. This result is stated in Thm. 6.1 and proved in App. C.2.
Theorem 6.1.
For every $\delta\in(0,1]$ and every value of the regularisation parameter $\lambda>0$, the RHS of (14) with $c=\sigma^{2}/\lambda$ and $\alpha=\lambda$ is strictly less than the UCB in Thm. 3.11 of (Abbasi-Yadkori,, 2012).
For every $\delta\in(0,1]$ and every value of the regularisation parameter $\eta>0$, the RHS of (14) with $c=\sigma^{2}/(1+\eta)$ and $\alpha=1+\eta$ is strictly less than the UCB in Thm. 2 of (Chowdhury and Gopalan,, 2017).
6.2 Cumulative Regret Bounds
We bound the cumulative regret in terms of two kernel-dependent quantities. The first is called the maximum information gain, and is defined as
$\gamma_{t}(\alpha):=\max_{x_{1},\dots,x_{t}}\left\{\tfrac{1}{2}\ln\left(\det\left(\tfrac{1}{\alpha}\boldsymbol{K}_{t}+\boldsymbol{I}\right)\right)\right\}.$
The second is defined as $\tau_{t}(\alpha):=|\mathcal{T}_{t}(\alpha)|$, where
$\mathcal{T}_{t}(\alpha)=\{s\in[t]:\tfrac{1}{\alpha}\rho_{\alpha,s-1}^{2}(x_{s})\geq 1\}.$ (16)
$\tau_{t}(\alpha)$ counts the number of rounds in which the scaled variance $\rho_{\alpha,s-1}^{2}(x_{s})/\alpha$, also known as the elliptical potential, is large. We call $\tau_{t}(\alpha)$ the elliptical potential count. First, we upper bound our radius quantity $\widetilde{R}_{\alpha,t}$ in terms of the maximum information gain. We show that when $\alpha=\sigma^{2}/c$, the quadratic terms in $\widetilde{R}_{\alpha,t}$ cancel out (see Lemma C.2), and we are left with
$\widetilde{R}_{\sigma^{2}/c,t}$ $\leq\sqrt{2\sigma^{2}\gamma_{t}(\sigma^{2}/c)+\tfrac{\sigma^{2}B^{2}}{c}+2\sigma^{2}\ln(\tfrac{1}{\delta})}.$ (17)
We now proceed to upper bound the cumulative regret of Kernel AMM-UCB with $\alpha=\sigma^{2}/c$. Using a standard optimism argument (see e.g. Prop. 1 in Russo and Van Roy, 2013 and also Lemma D.12 in App. D.3), the regret of any UCB algorithm at round $t$ is (with high probability) upper bounded by the difference between the upper and lower confidence bounds (at $x_{t}$) that it uses. For Kernel AMM-UCB, this bound is
$r_{t}\leq\tfrac{2}{\sqrt{\sigma^{2}/c}}\widetilde{R}_{\sigma^{2}/c,t-1}\rho_{\sigma^{2}/c,t-1}(x_{t}).$ (18)
Because the confidence bounds used by Kernel CMM-UCB (with the same $c$) and Kernel DMM-UCB (with the same $c$ and with $\alpha=\sigma^{2}/c\in A$) are never looser than those used by Kernel AMM-UCB, these algorithms also satisfy Eq. (18). Next, as is now somewhat common in analyses of (generalised) linear bandit algorithms (Gales et al.,, 2022; Lee et al.,, 2024; Janz et al.,, 2024), we split $r_{1:T}$ into $\sum_{t\in\mathcal{T}_{T}(\sigma^{2}/c)}r_{t}$ and $\sum_{t\notin\mathcal{T}_{T}(\sigma^{2}/c)}r_{t}$. Using $\sup_{x}|f^{\star}(x)|\leq C$, the first part can be bounded by $2C\tau_{T}(\sigma^{2}/c)$. Eq. (17), Eq. (18) and the Elliptical Potential Lemma (Lemma D.4) are used to bound the second part in terms of $\gamma_{T}(\sigma^{2}/c)$, which yields the regret bound in Thm 6.2 (see App. D.3 for a full proof).
Theorem 6.2.
$r_{1:T}\leq 2C\tau_{T}(\tfrac{\sigma^{2}}{c})+\sigma\sqrt{24T\gamma_{T}(\tfrac{\sigma^{2}}{c})(\gamma_{T}(\tfrac{\sigma^{2}}{c})+\tfrac{B^{2}}{2c}+\ln\tfrac{1}{\delta})}.$
Unless $\sigma$ is close enough to 00, the dominant term in this regret bound is the second term. Its dependence on $T$ and the maximum information gain matches that of the best existing regret bound for KernelUCB (Whitehouse et al.,, 2023), although the dependence on $\gamma_{T}$ is sub-optimal by a factor of $\sqrt{\gamma_{T}}$. However, due to the lower bound in (Lattimore,, 2023), this cannot be improved by using a better confidence sequence for $f^{\star}$.
We use bounds on $\gamma_{T}(\alpha)$ to derive new upper bounds $\tau_{T}(\alpha)$. Bounds on $\tau_{T}(\alpha)$ exist for the linear setting (e.g. Gales et al.,, 2022), where $k(x,x^{\prime})=x^{\top}x^{\prime}$, and $\gamma_{T}(\alpha)$ admits a relatively simple upper bound. However, for more general kernels, we must resort to more elaborate bounds on $\gamma_{T}(\alpha)$, which makes bounding $\tau_{T}(\alpha)$ rather complicated (cf. Lemma D.10 and Lemma D.11). These bounds allow us to: a) make the dependence of Thm. (6.2) on $T$ explicit for specific kernels; b) provide sub-$\sqrt{T}$ regret bounds for the noiseless case (Sec. 6.3).
For the RBF kernel, and with $c\propto 1$, we show in Prop. D.14 that the regret bound in Thm. 6.2 is $\mathcal{O}(\sqrt{T}(\ln(T))^{d+1})$. For the Matérn kernel with smoothness parameter $\nu$ and $c\propto T^{-\frac{d}{2d+2\nu}}$, we show in Prop. D.13 that the regret bound in Thm. 6.2 is $\mathcal{O}(T^{\frac{2d+\nu}{2d+2\nu}}(\ln(T))^{\frac{2\nu}{d+2\nu}})$. The same rates are obtained in (Whitehouse et al.,, 2023), which are the best known rates for (vanilla) KernelUCB.
6.3 The Noiseless Setting
In a 2022 COLT open problem, Vakili, (2022) discussed the lack of order-optimal regret bounds for noiseless kernel bandits. Though the open problem has since been resolved (Salgia et al.,, 2024), we believe it is still not known whether KernelUCB can achieve sub-$\sqrt{T}$ regret bounds in the noiseless setting (see Sec. 4.1. in Vakili,, 2022). We show that this is in fact the case. In the noiseless setting, the constraint $\|\boldsymbol{f}_{t}-\boldsymbol{y}_{t}\|_{2}\leq R_{t}$ in Lem. 4.2 can be replaced with $\|\boldsymbol{f}_{t}-\boldsymbol{y}_{t}\|_{2}=0$. Since $R_{t}=0$, the radius of the analytic confidence bound in Cor. 4.6 becomes
$\widetilde{R}_{\alpha,t}^{2}=\alpha B^{2}-\boldsymbol{y}_{t}^{\top}(\tfrac{1}{\alpha}\boldsymbol{K}_{t}+\boldsymbol{I})^{-1}\boldsymbol{y}_{t}\leq\alpha B^{2}$
Following the proof of Thm. 6.2, one can show that, for all $\alpha>0$ and $T\geq 1$,
$r_{1:T}\leq 2C\tau_{T}(\alpha)+\sqrt{12\alpha B^{2}T\gamma_{T}(\alpha)}.$
@@ 6.3 The Noiseless Setting
From here, in App. D.5, we show that for the RBF kernel, if $\alpha\propto 1/T$, then this bound becomes $r_{1:T}=\mathcal{O}(\ln(T)^{d+1})$. For the Matérn kernel, if $\alpha\propto T^{-\frac{2\nu^{2}+2\nu d}{2\nu^{2}+2\nu d+d^{2}}}$, then the regret bound becomes $r_{1:T}=\mathcal{O}(T^{\frac{\nu d+d^{2}}{2\nu^{2}+2\nu d+d^{2}}}\ln(T))$. Whenever $\nu>d/\sqrt{2}$, this bound grows slower that $\sqrt{T}$.
@@ 7 EXPERIMENTS
We aim to verify that when our confidence bounds replace existing ones, KernelUCB shows better empirical performance at a similar computational cost. We compare Kernel CMM-UCB, Kernel DMM-UCB, Kernel AMM-UCB, AY-GP-UCB, IGP-UCB and Random, which chooses actions uniformly at random.
We run each algorithm in several kernel bandit problems. To generate each reward function, we sample 20 inducing points $z_{1},\dots,z_{20}$ uniformly at random from the hypercube $[0,1]^{d}$ and a 20-dimensional weight vector $\boldsymbol{w}\in\mathbb{R}^{20}$ from a standard normal distribution. The test function is $f^{\star}(x)=b\sum_{i=1}^{20}w_{i}k(x,z_{i})$, with $b$ being set such that $\left\|f^{\star}\right\|_{\mathcal{H}}=B$. The noise variables $\epsilon_{1},\epsilon_{2},\dots$ are drawn independently from a normal distribution with mean 0 and standard deviation $\sigma$. In each round $t$, the action set $\mathcal{X}_{t}$ consists of 100 $d$-dimensional vectors drawn uniformly at random from the hypercube $[0,1]^{d}$. We present results for $d\in\{2,3,4\}$ and with the RBF and Matérn kernels with length scale $\ell\in\{0.2,0.5\}$.
We run each algorithm for $T=1000$ rounds with $\delta=0.01$ and the true values $B=10$ and $\sigma=0.1$. We run IGP-UCB with the recommend value of $\eta=2/T$. For each of our methods, we set $c=1$ for the RBF kernel and $c=T^{-\frac{d}{2d+2\nu}}$ for the Matérn kernel. Equivalently, for AY-GP-UCB, we set $\lambda=\sigma^{2}$ for the RBF kernel and $\lambda=\sigma^{2}T^{\frac{d}{2d+2\nu}}$ for the Matérn kernel. We run Kernel AMM-UCB with $\alpha=\sigma^{2}/c$ and Kernel DMM-UCB with $A=\{0.1\sigma^{2}/c,0.3\sigma^{2}/c,\sigma^{2}/c,3\sigma^{2}/c,10\sigma^{2}/c\}$. Due to prohibitive cost, we only run Kernel CMM-UCB for 200 rounds. All experiments were run on a single computer with an Intel i5-1145G7 CPU.
Fig. 2 shows the cumulative regret of each algorithm with $d=3$ (results with $d\in\{2,4\}$ are shown in App. F.2). The cumulative regret at $T=1000$ is displayed in Tab. 1. Our Kernel CMM-UCB algorithm consistently achieves the lowest cumulative regret over the first 200 steps, and is closely followed by Kernel DMM-UCB which achieves the lowest cumulative regret over 1000 steps. Fig. 4 (in App. F.1) shows the wall-clock time per step for each method. IGP-UCB, AY-GP-UCB and Kernel AMM-UCB take approximately the same time per step. Since we used a grid $A$ of size 5, the time per step for Kernel DMM-UCB is a little under 5 times that of Kernel AMM-UCB. The time per step for Kernel CMM-UCB grows at the fastest rate in $t$. At $t=200$, the time per step for Kernel CMM-UCB is more than 15 seconds, whereas the time per step for every other method is below 0.02 seconds.
@@ 8 CONCLUSION
In this paper, we developed new confidence bounds for sequential kernel regression. In our theoretical analysis, we proved that our new confidence bounds are always tighter than existing confidence bounds, and we showed that our KernelUCB algorithms can have sub-$\sqrt{T}$ cumulative regret in the noiseless setting. Our experimental results suggest that our Kernel DMM-UCB algorithm is the preferred choice. It closely matches the empirical performance of our effective but expensive Kernel CMM-UCB algorithm, yet it has roughly the same cost as our Kernel AMM-UCB method.
Though we applied our confidence bounds only to the kernel bandit problem, and the KernelUCB algorithm in particular, they are a generic tool that could be used anywhere that existing confidence bounds are used. For instance, one could use our confidence bounds in the $\pi$-GP-UCB algorithm (Janz et al.,, 2020), which would give a better rate for the Matérn kernel. In addition, one could apply our confidence bounds to kernelised reinforcement learning (Yang et al.,, 2020; Liu and Su,, 2022; Vakili and Olkhovskaya,, 2023) or adaptive control (Kakade et al.,, 2020).
Finally, we comment on some limitations of our work. We assume that the kernel as well as reasonably good upper bounds on $\sigma$ and $B$ are known in advance, which is often not the case in practice. While using a loose upper bound on $\sigma$ appears to have little impact, we found that using a loose bound on $B$ can harm empirical performance (see Fig. 7). In future work, we would like to address this limitation by using model selection methods to estimate $\left\|f^{\star}\right\|_{\mathcal{H}}$ and choose the kernel.